Determine the structural formula for each
8 PPM c. CSH10O2 =0 3H 3H 2H 2H 5 PPM CSH1002 d. 3H 3H 3H 1H PPM 45)
Answer #1

Molecular formula = C5H1002 DOU = 26+2+N-H-X - 2 2x5+2-10 2 TH NMR º multiplicity = 2nI+I, I=%, n= No. of proton I ) Proton aB) from spectrum (II): O= и сн ED Proton a → 2X044 +1 = 1 (singlet) - at f= 2.14 ppm (3H) in proton b→ 2 x0x[+1=1 (singlet) d